### v1

@@ Abstract
While mobile devices provide ever more compute power, improvements in DRAM bandwidth are much slower. This is unfortunate for large language model (LLM) token generation, which is heavily memory-bound. Previous work has proposed to leverage natural dynamic activation sparsity in ReLU-activated LLMs to reduce effective DRAM bandwidth per token. However, more recent LLMs use SwiGLU instead of ReLU, which result in little inherent sparsity. While SwiGLU activations can be pruned based on magnitude, the resulting sparsity patterns are difficult to predict, rendering previous approaches ineffective. To circumvent this issue, our work introduces Dynamic Input Pruning (DIP): a predictor-free dynamic sparsification approach, which preserves accuracy with minimal fine-tuning. DIP can further use lightweight LoRA adapters to regain some performance lost during sparsification. Lastly, we describe a novel cache-aware masking strategy, which considers the cache state and activation magnitude to further increase cache hit rate, improving LLM token rate on mobile devices. DIP outperforms other methods in terms of accuracy, memory and throughput trade-offs across simulated hardware settings. On Phi-3-Medium, DIP achieves a 46% reduction in memory and 40% increase in throughput with $<$ 0.1 loss in perplexity. We plan to open source the code for HW simulator, methods, and experiments in this paper.
@@ 1 Introduction
Specialized neural processing unit (NPU) accelerators for AI have become standard in recent smartphone SoCs. These processors enable efficient neural network inference on-device, opening the way for powerful on-device AI applications. However, as shown in Figure 2, memory size has not improved at the same pace as processing power Jhang et al. (2021); Kwon et al. (2023); Gholami et al. (2024). This memory bottleneck is especially evident for LLMs, which grew to billions and trillions of parameters, and for which all weights must be accessed to generate each individual output token. For instance, a model with 14B parameters like Phi-3-Medium quantized at INT4 precision occupies approximately 7GB, which approaches the average smartphone total DRAM capacity Counterpoint (2021); Grossi & Lorenz (2024).
Figure 1: Overview of the Dynamic Input Pruning method in a Gated MLP. The Cache-Aware masking component outputs a column sparsity mask based on current input activations and weights currently cached in DRAM. The unpruned columns are retrieved from DRAM cache, if available, or loaded from Flash in case of cache misses.
But, in fact, the OS and active applications already a significant portion of the total DRAM capacity, typically leaving a few GBs for an on-device LLM.
@@
To relax the DRAM capacity pressure, it has been previously proposed to load a model directly from flash, instead of first copying it in DRAM Alizadeh et al. (2023); Xue et al. (2024). By leveraging dynamic sparsity and introducing DRAM caching the latency overhead of loading a model from flash can be mitigated significantly. Dynamic sparsity Mirzadeh et al. (2023); Lee et al. (2024); Akhauri et al. (2024) exploits the fact that all operations during token generation are matrix-vector multiplications, and the fact that in LLMs with ReLU activations, vectors contain many zero elements. For elements in the activation vectors that are zero, we can skip loading the corresponding weights in the following layer. If sparsity predictors such as those introduced in DejaVu (Liu et al., 2023) are used to predict which elements in the activation vectors will be zero, the weights that generate zero-valued activation elements can also be ignored in the preceding layer. Additionally, a DRAM cache is used to cache active parameters from previous tokens. When a DRAM cache is used, parameters only need to be loaded from flash in case of a cache miss, i.e., when an active parameter is not already in cache from previous tokens. These approaches allow even Mixtral8x7B to run on mobile devices at reasonable token generation rates Xue et al. (2024).
@@
The high natural sparsity rates of ReLU LLMs (up to 90%) have two implications: 1) for each token, only a few active parameters need to be loaded from Flash, and 2) active parameters from multiple previous tokens can be cached, leading to higher cache hit rate and, as a result, reduced flash access. SwiGLU LLMs, however, can be pruned dynamically only up to approximately 50% before model accuracy is significantly impacted (see Section 3). Hence, to reach reasonable token generation rates at lower sparsity levels, we introduce a cache-aware variant of dynamic input pruning. This variant takes into account the current state of the cache when deciding which parameters to use. By biasing the parameter selection towards the ones that are already in cache, we can increase cache-hit rate and hence LLM throughput, while minimally affecting model accuracy.
@@ 2 Background and Related Work
Most dynamic sparsity approaches inherently rely on the sparse activation pattern induced by the ReLU non-linearity, however, the vast majority of recent LLM architectures Jiang et al. (2023); Abdin et al. (2024); Dubey et al. (2024) use non-linearities that do not naturally induce many hard zeros (cf. Figure 3). While ReLU-fication, the process of replacing the original network activations with ReLU activations, has shown promising results (Mirzadeh et al., 2023; Song et al., 2024b), though on the other hand it requires a computationally demanding retraining procedure which may fail to replicate the original performance.
@@ 3 Dynamic Sparsity on Modern LLMs
Figure 5: Diagrams describing the sparsity schemes for GLU pruning, Gate pruning, Predictive pruning and Input pruning. Activations with larger magnitude are visualized in darker colors, while dark blue indicates active weights for the current input.
@@ 3.2 Strategies for Dynamic Sparsity
$\widetilde{\bm{W}}^{d}_{:,i}=\begin{cases}\bm{0}\ \ \ \ \ \ \ \ \ \ \text{if }i\notin\text{topK}\left(|\text{GLU}({\mathbf{x}})_{i}|\right)\\ \bm{W}^{d}_{:,i}\ \ \ \ \ \text{o.w.}\end{cases}.$ (4)
@@ 6.1 Experimental setup
In DIP, we empirically calibrate the proportion of active units for a target MLP sparsity as described in Appendix B.1. For DIP-CA, we tune the cache-aware re-weighting hyper-parameter on WikiText-2 validation set and fix it to $\gamma=0.2$ for all experiments. We include an ablation showing the impact of changes in $\gamma$. Following previous work (Hu et al., 2021; Lee et al., 2024), we train LoRA adapters with rank 32 for a total of 1000 iterations using a knowledge distillation loss to match the original (dense) model logits. The training procedure is consistent for both DIP and CATS models. The adapters are applied to all the up, down and gate matrices for DIP, while only up and down matrices are adapted for the CATS method.
@@ 6.2 Results
In Table 1 we report more extensive numerical results on four modern LLMs. We consider 50% MLP sparsity for each method, and report the perplexity on WikiText-2 as well as the 5-shot accuracy on the MMLU benchmark. Results for different operating points can be found in Table 4 and Table 4 provided in Appendix C.

### v2

@@ Abstract
While mobile devices provide ever more compute power, improvements in DRAM bandwidth are much slower. This is unfortunate for large language model (LLM) token generation, which is heavily memory-bound. Previous work has proposed to leverage natural dynamic activation sparsity in ReLU-activated LLMs to reduce effective DRAM bandwidth per token. However, more recent LLMs use SwiGLU instead of ReLU, which results in little inherent sparsity. While SwiGLU activations can be pruned based on magnitude, the resulting sparsity patterns are difficult to predict, rendering previous approaches ineffective. To circumvent this issue, our work introduces Dynamic Input Pruning (DIP): a predictor-free dynamic sparsification approach, which preserves accuracy with minimal fine-tuning. DIP can further use lightweight LoRA adapters to regain some performance lost during sparsification. Lastly, we describe a novel cache-aware masking strategy, which considers the cache state and activation magnitude to further increase cache hit rate, improving LLM token rate on mobile devices. DIP outperforms other methods in terms of accuracy, memory and throughput trade-offs across simulated hardware settings. On Phi-3-Medium, DIP achieves a 46% reduction in memory and 40% increase in throughput with $<$ 0.1 loss in perplexity when compared to streaming the dense model from Flash. The open source code for HW simulator, methods, and experiments in this paper is available at https://github.com/Qualcomm-AI-research/dynamic-sparsity.
@@ 1 Introduction
Specialized neural processing unit (NPU) accelerators for AI have become standard in recent smartphone SoCs. These processors enable efficient neural network inference on-device, opening the way for powerful on-device AI applications. However, as shown in Figure 2, memory size has not improved at the same pace as processing power Jhang et al. (2021); Kwon et al. (2023); Gholami et al. (2024). This memory bottleneck is especially evident for LLMs, which grew to billions and trillions of parameters, and for which all weights must be accessed to generate each individual output token.
Figure 1: Overview of Cache-Aware Dynamic Input Pruning for a vector-matrix multiplication. The Cache-Aware masking component (DIP-CA) outputs a column sparsity mask based on current input activations and weights currently cached in DRAM. The subset of required weights are retrieved from DRAM cache, if available, or loaded from Flash in case of cache misses.
For instance, a model with 14B parameters like Phi-3-Medium quantized at INT4 precision occupies approximately 7GB, which approaches the average smartphone total DRAM capacity Counterpoint (2021); Grossi & Lorenz (2024).
@@
But, in fact, the OS and active applications already occupy a significant portion of the total DRAM capacity, typically leaving a few GBs free for an on-device LLM.
To relax the DRAM capacity pressure, it has been previously proposed to load a model directly from flash, instead of first copying it in DRAM Alizadeh et al. (2023); Xue et al. (2024). By leveraging dynamic sparsity and introducing DRAM caching the latency overhead of loading a model from flash can be mitigated significantly. Dynamic sparsity Mirzadeh et al. (2023); Lee et al. (2024); Akhauri et al. (2024) exploits the fact that all operations during token generation are matrix-vector multiplications, and the fact that in LLMs with ReLU activations, vectors contain many zero elements. For elements in the activation vectors that are zero, we can skip loading the corresponding columns in the following layer. If sparsity predictors such as those introduced in DejaVu (Liu et al., 2023) are used to predict which elements in the activation vectors will be zero, the weights that generate zero-valued activation elements can also be ignored in the preceding layer. Additionally, a DRAM cache is used to cache active parameters from previous tokens. When a DRAM cache is used, parameters only need to be loaded from flash in case of a cache miss, i.e., when an active parameter is not already in cache from previous tokens. These approaches allow even Mixtral8x7B to run on mobile devices at reasonable token generation rates Xue et al. (2024).
@@
The high natural sparsity rates of ReLU LLMs (up to 90%) have two implications: 1) for each token, only a few active parameters need to be loaded from Flash, and 2) active parameters from multiple previous tokens can be cached, leading to higher cache hit rate and, as a result, reduced flash access. SwiGLU LLMs, however, can be pruned dynamically only up to approximately 50% before model accuracy is significantly impacted (see Section 3). Hence, to reach reasonable token generation rates at lower sparsity levels, we introduce a cache-aware variant of dynamic input pruning, which we visualize in Figure 1. This variant takes into account the current state of the cache when deciding which parameters to use. By biasing the parameter selection towards the ones that are already in cache, we can increase cache-hit rate and hence LLM throughput, while minimally affecting model accuracy.
@@ 2 Background and Related Work
Most dynamic sparsity approaches inherently rely on the sparse activation pattern induced by the ReLU non-linearity, however, the vast majority of recent LLM architectures333 All the open-source LLMs benchmarked in Wang et al. (2024) use SwiGLU non-linearities.Jiang et al. (2023); Abdin et al. (2024); Dubey et al. (2024) use non-linearities that do not naturally induce many hard zeros (cf. Figure 3). While ReLU-fication, the process of replacing the original network activations with ReLU activations, has shown promising results (Mirzadeh et al., 2023; Song et al., 2024b), though on the other hand it requires a computationally demanding retraining procedure which may fail to replicate the original performance (SparseLLM, 2023; Lee et al., 2024).
@@ 3 Dynamic Sparsity on Modern LLMs
Figure 5: Diagrams describing the sparsity schemes for GLU pruning, Gate pruning, Predictive pruning and Input pruning. Activations with larger magnitude are visualized in darker colors, while dark blue indicates weights used for computation for the current input.
@@ 3.2 Strategies for Dynamic Sparsity
$\widetilde{\bm{W}}^{d}_{:,i}=\begin{cases}\bm{0}\ \ \ \ \ \ \ \ \ \ \text{if }i\notin\text{topK}\left(|\text{GLU}({\mathbf{x}})|\right)\\ \bm{W}^{d}_{:,i}\ \ \ \ \ \text{o.w.}\end{cases}.$ (4)
@@ 6.1 Experimental setup
In DIP, we empirically calibrate the proportion of active units for a target MLP sparsity as described in Appendix B.1. For DIP-CA, we tune the cache-aware re-weighting hyper-parameter on WikiText-2 validation set and fix it to $\gamma=0.2$ for all experiments. We include an ablation showing the impact of changes in $\gamma$. Following previous work (Hu et al., 2021; Lee et al., 2024), we train LoRA adapters with rank 32 for a total of 1000 iterations using a knowledge distillation loss to match the original (dense) model logits. The training procedure is consistent for DIP and CATS models. The adapters are applied to all the up, down and gate matrices for DIP, while only up and down are adapted for CATS.
@@ 6.2 Results
In Table 1 we report more extensive numerical results on four modern LLMs. We consider 50% MLP sparsity for each method, and report the perplexity on WikiText-2 as well as the 5-shot accuracy on the MMLU benchmark. Results for different operating points can be found in Table 4 and Table 4 provided in Appendix C, which also includes evaluations on a broader range of tasks.
@@ Appendix C Additional Results
Table 5 compare the performance of the same models at 50% MLP sparsity on a wide range of tasks including common reasonong (ARC, BoolQ, HellaSwag, PIQA, Wingogrande), multi-lingual reasonong (MGSM) and language understanding (MMLU-Pro). The evaluation procedures follows the protocol described in the Language Model Evaluation Harness (Gao et al., 2024). MGSM and MMLU-Pro require generation of multiple tokens and the corresponding accuracy is evaluated by considering exact matches. The reported values of MGSM refer to the average performance in 11 languages.
These additional results reinforce the trend we observed for perplexity and MMLU in Table 1, with DIP outperforming unstructured SparseGPT, DejaVU and CATS on most tasks and architectures.
@@
Table 5: Accuracy at 50% MLP sparsity on various language reasoning and understanding tasks.
Accuracy $\uparrow$
ARC (easy) ARC (challenge) BoolQ HellaSwag PIQA Winogrande MGSM MMLU-Pro
Phi3Med
Dense 85.40 61.43 88.53 65.07 80.85 76.56 32.95 52.69
GLU pruning (oracle) 85.31 60.58 88.75 65.09 81.39 75.37 33.24 52.19
SparseGPT (unstructured) 84.05 55.38 88.93 60.83 78.78 75.22 26.80 44.13
DejaVu 83.33 54.52 87.74 59.90 77.86 74.82 20.11 37.02
CATS 80.22 54.78 86.18 60.73 74.48 66.22 13.60 38.50
DIP 81.06 56.14 88.69 62.33 76.77 72.22 33.24 48.01
Phi3Mini
Dense 81.99 53.75 85.17 59.01 80.41 73.48 32.87 44.46
GLU pruning (oracle) 82.24 55.12 84.68 59.15 79.71 73.64 31.85 43.42
SparseGPT (unstructured) 78.28 47.44 80.61 53.92 77.26 70.64 20.22 30.30
DejaVu 75.29 45.14 74.59 43.74 76.50 62.98 6.33 24.17
CATS 75.55 45.65 65.99 52.50 72.91 62.12 6.40 21.76
DIP 79.42 53.33 83.15 57.22 77.31 68.43 26.36 36.66
Llama8B
Dense 80.09 50.06 81.35 60.19 79.71 72.38 9.75 34.47
GLU prining (oracle) 79.38 49.83 80.52 60.21 79.82 73.24 10.07 33.69
SparseGPT (unstructured) 75.42 41.30 78.99 54.58 77.09 71.11 5.13 24.00
DejaVu 67.59 39.76 74.89 53.89 76.44 68.59 3.27 20.10
CATS 54.38 32.25 56.70 45.43 70.24 57.54 0.47 7.65
DIP 77.57 46.50 79.51 57.90 77.53 71.11 7.35 29.48
Mistral7B
Dense 80.89 50.43 83.61 61.21 80.58 73.88 4.40 30.39
GLU pruning (oracle) 80.51 50.26 83.33 61.59 80.63 73.09 4.47 30.18
SparseGPT (unstructured) 76.68 44.88 76.24 57.05 78.18 71.82 4.11 24.19
DejaVu 77.40 45.90 79.24 60.47 79.33 72.22 2.87 23.55
CATS 51.56 32.00 73.85 49.67 70.02 61.64 0.47 5.66
DIP 79.12 48.98 83.15 60.70 79.54 72.53 4.29 28.02
@@ Appendix D Ablation on hardware specifications
Appendix E Artifact Appendix
E.1 Abstract
This Artifact Appendix describes the experimental workflow, artifacts and results from this paper submitted for the Artifact Evaluation at MLSys 2025. The experiments described in this paper can be reproduced through the open source code released as a code artifact Federici et al. (2025). The code should be executed in a Python 3 environment as described in this appendix. The setup relies on GPU hardware to accelerate the experiments, with the smallest model requiring 40GB of VRAM. Datasets and models used in this paper can be obtained from Huggingface [^4] https://huggingface.co/models.
E.2 Artifact check-list (meta-information)
  • Algorithm: Efficient LLM inference.
  • Models: Phi-3-Medium, Phi-3-Mini, Llama-v3-8B, Mistral-v01-7B.
  • Data sets: Wikitext, MMLU, SlimPajama (optional).
  • Run-time environment: Python 3.8 and above interpreter and latest pip version to install the required packages.
  • Hardware: A torch-enabled GPU with 80GB of VRAM is required for most models. For Phi-3-Mini, 40GB of VRAM are enough.
  • Execution: python scripts/run_experiment.py –help
  • Metrics: Perplexity, 5-shot accuracy, throughput, memory footprint, MLP density.
  • Output: Metrics for a given method, model, dataset and hardware simulator setup.
  • Experiments: Experiments to reproduce Table 1 (perplexity and 5-shot accuracy at 50% sparsity) and Table 2 (throughput at given perplexity increase) in the paper.
  • How much disk space required (approximately)?: Less than 1 MB for the code, Up to 1GB for the pyhon environment setup. Up to 100GB for the LLM models.
  • How much time is needed to prepare workflow (approximately)?: A few minutes for the environment setup and download each model and dataset.
  • How much time is needed to complete experiments (approximately)?: Less than 10 minutes for a single run with Phi-3-Mini on Wikitext. More than 200 GPU hours to reproduce all experiments.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: BSD 3-Clause Clear License.
  • Data licenses (if publicly available)?: As specified for each Huggingface dataset.
  • Workflow framework used?: No.
  • Archived (provide DOI)?: 10.5281/zenodo.15088634
E.3 Description
E.3.1 How delivered
Our source code and instructions for installation, setup and experiment reproduction are publicly available at: https://github.com/Qualcomm-AI-research/dynamic-sparsity.
E.3.2 Hardware dependencies
We recommend testing on a Linux machine with torch-enabled GPU. We used a x86_64 architecture with an AMD 64-core processor, Nvidia H100 with 80GB of VRAM and CUDA 12.2.
E.3.3 Software dependencies
Python 3.8 and above interpreter and latest pip version to install the required packages. Tested with Python 3.8 and pip 24.3.1. Packages are listed in requirements.txt.
E.3.4 Data sets
The Wikitext and MMLU data sets are required to reproduce the main experiments, and SlimPajama is needed when running the method variation employing LoRA finetuning.
E.4 Installation
Python 3 environment with installed packages as listed in requirements.txt. Dataset and models must be downloaded from Huggingface. All details on installation and setup are provided in the Getting Started section of the README documentation.
E.5 Experiment workflow
In the Usage subsection of the README file we explain how to run experiments with the released code, including a description of all the main parameters regulating an experiment.
E.6 Evaluation and expected result
In the Reproducing Results section, we explain how to prepare and launch the experiments to reproduce the results in Table 1 (perplexity and 5-shot accuracy at 50% sparsity) and Table 2 (throughput at given perplexity increase) in the paper.
E.7 Experiment customization
The provided source code implements a set of dynamic sparsity methods for efficient LLM inference. We showcase in our paper experimental results on a selected set of data sets and models, but experimenting with different models and data sets is also possible with minimal code changes. Similarly, we release in our source code a hardware simulator to assess the system throughput under different configurations. These configurations are flexible and can be easily extended to simulate new scenarios. As an example, we included in /scripts/config/hw_simulator/processor the configurations for a variety of Apple processors based on the specifications provided in Wikipedia. An overview of the main components in our codebase is included in the Repository Structure section.